OpenStudy (anonymous):

2x – 3(x + 4) = - 5, then x =

OpenStudy (anonymous):

2x-3x-12=-5 ..............we multiplied 3 with (x+3) -x-12=-5 -x=12-5 -x=7.................x=-7